In 7 years, Maria will be four times as old as Angelica will be then. The sum of their ages now is 71. How old will each of them be in 5 years?
Found 2 solutions by ichudov, srirajdhaks:
Answer by ichudov(507)

All you have to do is express it as equations:
Maria + 7 = 4*(Angelica +7)
Maria + Angelica = 71
Replace Maria with M and Angelica with A
M=4*A = -7-4*7
M+A = 71
This is a classic linear system of equations, for which I have a solver that shows work.

Answer by srirajdhaks(1)

(Name) (Age now) (Age after 5 years)
(Maria) (7+4x) (7+4x+5 = 12+4x)
(Angelica) (7+x) (7+x+5 = 12+x)
The sum of their ages now is 71
12+4x+12+x =71
24+5x =71
5x =71-24
5x =47
x =47/5 =9.4
Maria : 7+4(9.4) =44.6
Angelica : 7+9.4 =16.4
Verify the answer:
12+4(9.4)+12+9.4 =71
71 =71